# date20082024

### v1

@@ Introduction
Fig. 4: Quench response of the driven emitters in the resolved vacuum Rabi splitting regime. a, Spin projection $J_{z}$ as a function of total drive duration $t$ accompanied by the theory predictions (solid lines). The data are taken by suddenly turning on the drive in 100100100~{}100ns and holding the drive constant for a variable amount of time, before extinguishing the drive and measuring the spin projection $J_{z}$. Traces with different colors represent different drive strengths with fixed atom number $N=8.0(8)\times 10^{3}$. The traces with $\Omega_{d}/\Omega_{c}=0.57,0.74$ exhibit characteristic oscillations associated with the vacuum Rabi frequency $2g\sqrt{N}$. The trace with $\Omega_{d}/\Omega_{c}=1.0$ should normally relax to $J_{z}=0$ as it is at the critical point of the CRF superradiant transition, but the quench causes $J_{z}$ quickly to shoot above 00 in $1~{}\mu$s, after which it remains in the normal phase. b, Spin projection $J_{z}$ as a function of $\Omega_{d}/\Omega_{c}$ at the end of 9.3 $\mu$s drive for quenching on the drive (orange) and ramping on the drive (purple) with $N=8.3(3)\times 10^{3}$. The procedure for ramping the drive is the same as that for Fig. 2. Under the same drive strength, the spin projection $J_{z}$ for the quenched experiments agrees with the full numerical simulation (orange solid line). However, they are consistently above or equal to $J_{z}$ obtained by ramping the drive, suggesting that with a quench the system can overshoot and fail to relax to the true steady state predicted by the mean-field theory (purple solid line) on experimental timescales.
@@
Fig. 5: Invariance of the critical point to detuning $\Delta_{ca}$ of the cavity from atomic resonance. We show the measured spin projection $J_{z}$ as a function of normalized Rabi frequency $\Omega_{d}/\Omega_{c}$ with $\Omega_{d}$ being the Rabi frequency that would be established in the bare cavity when $\Delta_{ca}=0$, and critical Rabi frequency $\Omega_{c}=0.35NC\gamma$. The different color traces are taken at different cavity-atom detunings $\Delta_{ca}$ spanning many bare cavity linewidths $-5\kappa$ to $+5\kappa$ and with fixed atom number, while the drive is kept on resonance with the atoms ($\omega_{d}=\omega_{a}$). Despite such large changes, the traces all closely collapse on top of each other below the transition, showing the predicted insensitivity to the cavity detuning. Full numerical simulation suggests the dependence on detuning on $\Delta_{ca}$ is small despite the presence of the inhomogeneous coupling (see SI), so for clarity here we are only showing the black solid line for the case of $\Delta_{ca}=0$.
@@
While the ramp employed in Fig. 2 is designed to be slow enough to prepare the system close to the superradiant steady state, if the system far away from this steady state suddenly sees a quenched drive, the atoms would not have enough time to generate a field to cancel the external drive and may start to oscillate. Hence, the resolved VRS regime features an additional region of bistability compared to the non-resolved VRS regime [48] where the short-time dynamical behaviour of the system, either settling into the CRF steady state or oscillating persistently, is determined by the initial conditions. We can characterize this effect by comparing the $J_{z}$ values obtained after driving the system for $9.3~{}\mu$s following a quench or ramp of the drive. Fig. 4b shows, consistently with theory predictions, that the post-quench $J_{z}$ is always above the post-ramp $J_{z}$, and the difference increases as we approach the critical point $\Omega_{d}/\Omega_{c}=1$. Moreover, for $\Omega_{d}/\Omega_{c}$ slightly smaller than 1, we observe that the system appears to enter the normal phase after a quench even though it remains in the superradiant phase after a ramp. This highlights how, in the resolved VRS regime, our ability to access the superradiant phase can be influenced by the dynamical response to the drive.
@@
In summary, we have observed the superradiant phase transition in the cooperative resonance fluorescence model predicted more than 40 years ago. We also witnessed how spontaneous emission melted the continuous transition into a first order transition. In the future, it will be interesting to explore the properties of the steady states in the presence of additional Hamiltonian interactions [61] or at the level of quantum fluctuations since the spins get more squeezed as the drive strength approaches the critical point from below [17, 20]. Alternatively, by introducing a third atomic level to our system, one can also potentially generate squeezed states between the ground state and the third state via the generation of a dissipative Berry phase [21]. These schemes are all compatible with ultra-long-lived clock transitions in alkali-earth atoms and thus can be utilized to improve state-of-the-art atomic clocks [74, 75]. Furthermore, carefully monitoring the light leaking out of the cavity, in combination with post-selection and feedback operations, opens up the possibility of implementing exotic out-of-equilibrium phenomena such as measurement-induced phase transitions [76, 50].
@@ Acknowledgments
    Numerical simulations that incorporate the short time effects of spontaneous emission indicate that spontaneous emission further shifts the transition point to $\Omega_{c}^{nh,*}\equiv 0.78\Omega_{c}^{h}$ [see Fig S1(a)]. Inhomogeneous broadening of the atomic transition further shifts the transition point to $\Omega_{c}\equiv 0.70\Omega_{c}^{h}$, which is what is observed experimentally. Furthermore, once we rescale each curve by their respective critical point, we find that they fall on top of each other [Fig. S1(b)].
@@
Fig. S2: (a) We show the long time steady state weighted inversion $\tilde{J}_{z}$ as a function of $\Omega_{d}/\Omega_{c}^{h}$ for three different scenarios: homogeneous couplings and no broadening (green), inhomogeneous couplings and no broadening (red), and inhomogeneous couplings with broadening (purple, $\delta_{\text{max}}=2\pi\times 125$ kHz). All curves incorporate the effects of spontaneous emission. (b) Comparison between the steady state profile for $\delta_{\text{max}}=2\pi\times 125$ kHz and time evolution for $T_{\text{hold}}=100\mu\text{s},\,300\mu\text{s}$.

### v2

@@ Introduction
Fig. 4: Quench response of driven emitters in the resolved vacuum Rabi splitting regime. a, The spin projection $J_{z}$ as a function of the total drive duration $t$ accompanied by theory predictions (solid curves). The data are taken by suddenly turning on the drive in 100100100~{}100ns and holding the drive constant for a variable amount of time, before extinguishing the drive and measuring the spin projection $J_{z}$. Traces with different colors represent different drive strengths with fixed atom number $N=8.0(8)\times 10^{3}$. The traces with $\Omega_{d}/\Omega_{c}=0.57,0.74$ exhibit characteristic oscillations associated with the vacuum Rabi frequency $2g\sqrt{N}$. The trace with $\Omega_{d}/\Omega_{c}=1.0$ should normally relax to $J_{z}=0$ as it is at the critical point of the CRF superradiant transition, but the quench causes $J_{z}$ quickly to shoot above 00 in $1~{}\mu$s, after which it remains in the normal phase. b, The spin projection $J_{z}$ as a function of $\Omega_{d}/\Omega_{c}$ at the end of 9.3 $\mu$s drive for quenching (orange) and ramping on the drive (purple) with $N=8.3(3)\times 10^{3}$. The ramping procedure is the same as that for Fig. 2. Both datasets agree well with numerical simulations (solid curves). However, for the same $\Omega_{d}$, the measured $J_{z}$ in quenched experiments is consistently above the $J_{z}$ obtained in ramped experiments. This suggests that with a quench the system can overshoot and fail to relax to the true steady state predicted by the mean-field theory (purple solid line) on experimental timescales.
@@
Fig. 5: Invariance of the critical point to detuning $\Delta_{ca}$ of the cavity from atomic resonance. We show the measured spin projection $J_{z}$ as a function of $\Omega_{d}$, defined as the Rabi frequency that would be established in the bare cavity when $\Delta_{ca}=0$, normalized to the critical Rabi frequency $\Omega_{c}=0.35NC\gamma$. The different colours of data are taken at different cavity-atom detunings $\Delta_{ca}$ spanning many bare cavity linewidths (from $-5\kappa$ to $+5\kappa$) and with fixed atom number, while the drive is kept on resonance with the atoms ($\omega_{d}=\omega_{a}$). Despite such large changes, the datasets all overlap with each other below the transition, showing the predicted insensitivity to the cavity detuning. Full numerical simulations suggest the dependence on detuning on $\Delta_{ca}$ is small despite the presence of the inhomogeneous coupling (see SI), so for clarity here we only show simulations for $\Delta_{ca}=0$ (black curve).
@@
While the ramp employed in Fig. 2 is designed to be slow enough to prepare the system close to the superradiant steady state, if the system far away from this steady state is suddenly subject to a quenched drive, the atoms would not have enough time to generate a field to cancel the external drive and may start to oscillate. Hence, the resolved VRS regime features an additional region of bistability compared to the non-resolved VRS regime [48] where the short-time dynamical behaviour of the system, either settling into the CRF steady state or oscillating persistently, is determined by the initial conditions. We can characterize this effect by comparing the $J_{z}$ values obtained after driving the system for $9.3~{}\mu$s following a quench or ramp of the drive. Fig. 4b shows, consistently with theory predictions, that the post-quench $J_{z}$ is always above the post-ramp $J_{z}$, and the difference increases as we approach the critical point $\Omega_{d}/\Omega_{c}=1$. Moreover, for $\Omega_{d}/\Omega_{c}$ slightly smaller than 1, we observe that the system appears to enter the normal phase after a quench even though it remains in the superradiant phase after a ramp. This highlights how, in the resolved VRS regime, our ability to access the superradiant phase can be influenced by the dynamical response to the drive.
@@
In summary, we have observed the superradiant phase transition in the cooperative resonance fluorescence model predicted more than 40 years ago. We also witnessed how spontaneous emission melted the continuous transition into a first order transition. In the future, it will be interesting to explore the properties of the steady states in the presence of additional Hamiltonian interactions [61] or at the level of quantum fluctuations since the spins get more squeezed as the drive strength approaches the critical point from below [17, 20]. Alternatively, by introducing a third atomic level to our system, one can also potentially generate squeezed states between the ground state and the third state via the generation of a dissipative Berry phase [21]. These schemes are all compatible with ultra-long-lived clock transitions in alkaline earth atoms and thus can be utilized to improve state-of-the-art atomic clocks [74, 75]. Furthermore, carefully monitoring the light leaking out of the cavity, in combination with post-selection and feedback operations, opens up the possibility of implementing exotic out-of-equilibrium phenomena such as measurement-induced phase transitions [76, 50].
@@ Acknowledgments
    Numerical simulations that incorporate the short time effects of spontaneous emission indicate that spontaneous emission further shifts the transition point to $\Omega_{c}^{nh,*}\equiv 0.78\Omega_{c}^{h}$ [see Fig. S1(a)]. Inhomogeneous broadening of the atomic transition further shifts the transition point to $\Omega_{c}\equiv 0.70\Omega_{c}^{h}$, which is what is observed experimentally. Furthermore, once we rescale each curve by their respective critical point, we find that they fall on top of each other [Fig. S1(b)].
@@
Fig. S2: (a) We show the long-time steady state weighted inversion $\tilde{J}_{z}$ as a function of $\Omega_{d}/\Omega_{c}^{h}$ for three different scenarios: homogeneous couplings and no broadening (green), inhomogeneous couplings and no broadening (red), and inhomogeneous couplings with broadening (purple, $\delta_{\text{max}}=2\pi\times 125$ kHz). All curves incorporate the effects of spontaneous emission. (b) Comparison between the steady state profile for $\delta_{\text{max}}=2\pi\times 125$ kHz and time evolution for $T_{\text{hold}}=100\mu\text{s},\,300\mu\text{s}$.